# PGOT: A Physics-Geometry Operator Transformer for Complex PDEs

### v1

@@ 1 Introduction
Partial Differential Equations (PDEs), derived from fundamental principles such as conservation of mass and energy, are ubiquitous in modeling complex phenomena across science and engineering, including fluid dynamics, structural mechanics, and climate modeling (Cao et al., 2025). Traditional numerical methods, such as the Finite Element Method (FEM) (Zienkiewicz & Taylor, 2005) and Finite Volume Method (FVM) (Eymard et al., 2000), approximate solutions by discretizing PDEs onto meshes (Chen et al., 2025). However, their reliance on dense mesh discretization imposes a heavy computational burden, often requiring days or even months for multi-physics problems necessitating fine spatiotemporal resolution. Recently, Neural PDE Solvers have emerged as a promising paradigm, capable of learning resolution-invariant solution mappings between infinite-dimensional function spaces (Wang et al., 2025a; Li et al., 2025). By training on offline data, these operators can predict families of solutions under varying parameters orders of magnitude faster than traditional solvers, offering a promising direction for real-time scientific computing.
@@
Early spectral-based neural operators, such as the Fourier Neural Operator (FNO) (Li et al., 2021), rely on the fast fourier transform. While they possess the remarkable ability to learn resolution-independent operators, they are inherently limited to regular grids, restricting their applicability to complex geometries and unstructured meshes. To address this, graph-based neural operators (Hao et al., 2023; Li et al., 2023b) treat input and output functions as graphs, employing message-passing mechanisms to generalize across different domain geometries. However, graph methods are often constrained by local receptive fields, making it difficult to effectively capture global long-range dependencies and suffering from low efficiency on large-scale nodes.
This necessitates the adoption of Transformers. The self-attention mechanism in Transformers is mathematically analogous to learning a global Green’s function (Cao, 2021), making it ideally suited for modeling non-local physical interactions (Alkin et al., 2024; Koshizuka et al., 2024). However, a fundamental scalability barrier remains: the standard self-attention mechanism exhibits quadratic computational complexity $\mathcal{O}(N^{2})$ with respect to the number of mesh points $N$. Given that scientific simulations typically involve $10^{5}$ to $10^{7}$ grid points, standard Transformers are computationally intractable for high-resolution physics. Consequently, the research frontier has shifted towards efficient linear transformers (Li et al., 2023c; Luo et al., 2025; Wu et al., 2024). Recent state-of-the-art methods employ strategies such as low-rank approximations or token clustering to reduce the effective sequence length, achieving linear complexity $\mathcal{O}(N)$ (Rao et al., 2021; Xiao et al., 2024). While these architectures successfully improve throughput, they often do so at the cost of geometric and physical fidelity, as follows:
@@ 2.1 Physics-Informed Neural Networks
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) represent a foundational paradigm for solving PDEs by embedding governing equations directly into the loss function. PINNs parameterize the solution as a neural network and minimize the PDE residual at collocation points, enabling mesh-free training without labeled simulation data. However, PINNs face fundamental limitations for the problems we target. First, PINNs solve individual PDE instances, requiring complete retraining for each new parameter configuration or boundary condition—a prohibitive cost for design optimization requiring thousands of evaluations. Second, for complex geometries and multi-physics coupling, PINNs frequently encounter ill-conditioning issues (Zhang et al., 2025), often leading to convergence failures or accuracy degradation. Although recent research has alleviated these problems through parameterization (Cho et al., 2024), meta-learning (Cheng & Alkhalifah, 2025), and transfer learning (Wang et al., 2025b) strategies, PINNs still struggle to maintain accuracy and efficiency when the configuration space becomes sufficiently large. These limitations motivate the development of the neural operator paradigm.
@@ 2.2 Neural Operators for PDEs
Neural operators have emerged as a powerful paradigm for learning resolution-invariant mappings between infinite-dimensional function spaces (Kovachki et al., 2023). DeepONet (Lu et al., 2021) pioneered this field by approximating operators through a Branch-Trunk architecture. Subsequently, the Fourier Neural Operator (FNO) (Li et al., 2021) established new benchmarks by parameterizing the integral kernel directly in the frequency domain using FFT. To address the limitations of standard FNO, numerous variants have been proposed: U-NO (Rahman et al., 2023) and U-FNO (Wen et al., 2022) integrate U-Net structures to capture multi-scale features, while F-FNO (Tran et al., 2023) improves computational efficiency by factorizing the Fourier representation across dimensions. To handle irregular geometries beyond rectangular grids, Geo-FNO (Li et al., 2023a) maps physical domains to latent uniform grids, while LSM (Wu et al., 2023) and LNO (Wang & Wang, 2024) apply spectral methods within learned low-dimensional latent spaces to tackle high-dimensional PDE problems. Furthermore, to mitigate the global nature of Fourier transforms that typically leads to loss of local details, wavelet-based methods such as WMT (Gupta et al., 2021) and WNO (Tripura & Chakraborty, 2023) have been introduced to leverage the time-frequency localization properties of wavelets for capturing local variations. For modeling complex unstructured meshes, geometric deep learning offers an alternative perspective. Graph Neural Operators (GNO) (Li et al., 2020) and MeshGraphNet (Pfaff et al., 2021) treat physical domains as graphs, employing message-passing mechanisms to handle complex boundaries. To combine global spectral processing capabilities with local geometric flexibility, hybrid architectures have emerged. For instance, GINO (Li et al., 2023b) integrates Geo-FNO’s global spectral capabilities with GNO’s local graph processing, while 3D-GeoCA (Deng et al., 2024) further incorporates pre-trained 3D vision backbones to enhance geometric feature extraction. However, graph-based methods typically suffer from geometric instability (Morris et al., 2023; Luo et al., 2025) and incur high computational overhead when scaling to million-scale mesh points due to neighborhood search and message passing.
@@ 2.3 Transformers for PDEs
The Transformer architecture has garnered significant attention in scientific computing due to its global receptive field and superior modeling capabilities. Early works such as OFormer (Li et al., 2023c) and HT-Net (Liu et al., 2024) demonstrated the potential of attention mechanisms in capturing non-local physical interactions. However, the quadratic complexity $\mathcal{O}(N^{2})$ of standard attention remains a bottleneck restricting its scalability. To achieve linear complexity $\mathcal{O}(N)$, recent research has explored diverse pathways: AFNO (Guibas et al., 2022), FNet  (Lee-Thorp et al., 2022), and GFNet (Rao et al., 2021) attempt token mixing in the Fourier domain, while SAOT (Zhou et al., 2025) further integrates wavelet attention to mitigate aliasing issues; simultaneously, Galerkin Transformer (Cao, 2021), GNOT (Hao et al., 2023), and ONO (Xiao et al., 2024) utilize kernel tricks or low-rank approximations to achieve linearization; whereas methods represented by FactFormer  (Li et al., 2023d) and Transolver (Wu et al., 2024) employ axial decomposition or token aggregation strategies, respectively. Despite their significant efficiency improvements, these dimensionality reduction or aggregation operations act as implicit spatial low-pass filters, inevitably discarding high-frequency geometric details. Furthermore, they typically allocate a uniform computational budget across the entire domain, failing to adapt to physical heterogeneity in flow fields.
@@ 4.1 Experimental Setup
Our evaluation encompasses six benchmarks with varying geometric complexities. The standard benchmarks (Li et al., 2021, 2023a) include Elasticity (inner stress prediction on 2D point clouds), Plasticity (temporal displacement forecasting under external forces), Airfoil (transonic flow simulation with shock waves), and Pipe (velocity field estimation in curved structures). For industrial-scale validation, we employ Shape-Net Car (Umetani & Bickel, 2018) for automotive aerodynamic simulation on hybrid 3D geometries, and AirfRANS (Bonnet et al., 2022) for high-fidelity RANS simulation on NACA airfoils under varying Reynolds numbers and angles of attack. Detailed dataset descriptions are provided in Appendix A.1.
@@
We compare against representative methods from three categories: spectral neural operators (FNO (Li et al., 2021), Geo-FNO (Li et al., 2023a), U-NO (Rahman et al., 2023), F-FNO (Tran et al., 2023), LSM (Wu et al., 2023), WMT (Gupta et al., 2021)), Transformer-based approaches (Galerkin Transformer (Cao, 2021), OFormer (Li et al., 2023c), GNOT (Hao et al., 2023), ONO (Xiao et al., 2024), FactFormer (Li et al., 2023d), IPOT (Lee & Oh, 2024), Transolver (Wu et al., 2024), SAOT (Zhou et al., 2025)), and geometric deep learning methods (PointNet (Qi et al., 2017), Graph U-Net (Gao & Ji, 2019), GraphSAGE (Hamilton et al., 2017), MeshGraphNet (Pfaff et al., 2021), GNO (Li et al., 2020), GINO (Li et al., 2023b), 3D-GeoCA (Deng et al., 2024)).
@@
We adopt the relative $L^{2}$ error as the primary evaluation metric. For industrial benchmarks, we additionally report aerodynamic coefficient errors and Spearman’s rank correlation. PGOT is configured with $L=8$ PhysGeoBlocks, with other hyperparameters adapted to each task. All experiments are conducted on NVIDIA A100 GPUs and repeated three times. Comprehensive details on metrics, training configurations, and hyperparameters are provided in Appendix A.2.
@@ 4.2 Main Results
Table 2 summarizes results on large-scale industrial benchmarks. On Shape-Net Car, PGOT achieves the best performance across all metrics, with 9.1% reduction in volumetric field error and 10.6% improvement in drag coefficient prediction compared to Transolver. The Spearman correlation $\rho_{D}=0.9926$ indicates excellent ranking capability for design optimization. Figure 5(b) visualizes these results, showing excellent agreement between PGOT’s predicted streamlines and ground truth, with superior accuracy in geometrically complex regions like wheel wells. On AirfRANS, PGOT demonstrates substantial improvements: 32.4% reduction in volumetric field error, 81.3% reduction in surface field error, and 12.2% improvement in lift coefficient prediction. As shown in Figure 5(a), while Transolver and SAOT produce radial error patterns, PGOT achieves significantly cleaner predictions with errors primarily concentrated in the far-field region. These results underscore PGOT’s capability in handling complex unstructured meshes with over 32,000 points while maintaining high prediction fidelity.
@@ 4.3 Ablation Studies
Table 4 investigates the effect of multi-scale geometric encoding. Single-scale encoding ($S=1$) performs poorly across all benchmarks, confirming that aggregation-induced geometric aliasing significantly degrades prediction quality. Performance improves substantially and reaches optimality at $S=2$, demonstrating that a dual-scale representation is sufficient to capture both global topology and local boundary details. Interestingly, further increasing the scales to $S=3$ or $S=4$ leads to performance degradation. This suggests that excessive high-frequency geometric priors may introduce spectral noise or redundancy, which interferes with the model’s ability to learn robust physical dynamics.
@@ 5 Conclusion
Appendix A Implementation Details
A.1 Dataset Descriptions
@@
This industrial benchmark focuses on automotive aerodynamic simulation at 72 km/h driving conditions. The task requires simultaneous prediction of surface pressure and surrounding air velocity. Vehicle geometries are sourced from the “car” category of the ShapeNet repository and discretized into unstructured meshes with approximately 32,000 points combining surface and volumetric elements. The input is preprocessed to include mesh point positions, signed distance functions, and normal vectors. The predicted fields can be used to compute drag coefficients for design optimization. The training set contains 789 samples, and the test set contains 100 samples (Umetani & Bickel, 2018).
@@ A.2 Evaluation Metrics
A.3 Training and Model Configurations
Following the conventions established in prior work (Li et al., 2021, 2023a; Wu et al., 2024; Zhou et al., 2025), our model is trained using the relative $L^{2}$ loss. Table 8 summarizes the training configurations shared across all methods for fair comparison, and Table 9 presents PGOT-specific hyperparameters.
@@ A.3 Training and Model Configurations
For Shape-Net Car (Umetani & Bickel, 2018) and AirfRANS (Bonnet et al., 2022), $\mathcal{L}_{v}$ and $\mathcal{L}_{s}$ denote the relative $L^{2}$ losses on volumetric and surface fields.
@@ A.4 Baseline Implementations
For FNO (Li et al., 2021), Geo-FNO (Li et al., 2023a), WMT (Gupta et al., 2021), U-FNO (Wen et al., 2022), U-NO (Rahman et al., 2023), F-FNO (Tran et al., 2023), and LSM (Wu et al., 2023), we report results from the original publications (Wu et al., 2024; Zhou et al., 2025; Li et al., 2021).
@@
For Galerkin Transformer (Cao, 2021), OFormer (Li et al., 2023c), GNOT (Hao et al., 2023), HT-Net (Liu et al., 2024), FactFormer (Li et al., 2023d), and ONO, we report results from the original publications (Wu et al., 2024; Zhou et al., 2025; Li et al., 2021). For Transolver (Wu et al., 2024) and SAOT (Zhou et al., 2025), we reproduce results based on their official codebases with all hyperparameter configurations kept consistent with PGOT to ensure fair comparison.
@@
For MLP, GraphSAGE (Hamilton et al., 2017), PointNet (Qi et al., 2017), Graph U-Net (Gao & Ji, 2019), MeshGraphNet (Pfaff et al., 2021), GNO (Li et al., 2020), GINO (Li et al., 2023b), and 3D-GeoCA (Deng et al., 2024), we report results from the original publications.
Appendix B Supplementary Results
B.1 Ablation Results

### v2

@@ 1 Introduction
Partial Differential Equations (PDEs), derived from fundamental principles such as conservation of mass and energy, are ubiquitous in modeling complex phenomena across science and engineering, including fluid dynamics, structural mechanics, and climate modeling (Cao et al., 2025). Traditional numerical methods, such as the Finite Element Method (FEM) (Zienkiewicz and Taylor, 2005) and Finite Volume Method (FVM) (Eymard et al., 2000), approximate solutions by discretizing PDEs onto meshes (Chen and others, 2025; Zhang et al., 2024). However, their reliance on dense mesh discretization imposes a heavy computational burden, often requiring much time and effort for multi-physics problems necessitating fine spatiotemporal resolution. Recently, Neural PDE Solvers have emerged as a promising paradigm, capable of learning resolution-invariant solution mappings between infinite-dimensional function spaces (Wang et al., 2025a; Li et al., 2025). By training on offline data, these operators can predict families of solutions under varying parameters orders of magnitude faster than traditional solvers, offering a promising direction for real-time scientific computing.
@@
Early spectral-based neural operators, such as the Fourier Neural Operator (FNO) (Li et al., 2021), rely on the fast fourier transform. While they possess the remarkable ability to learn resolution-independent operators, they are inherently limited to regular grids, restricting their applicability to complex geometries and unstructured meshes. To address this, graph-based neural operators (Hao et al., 2023; Li et al., 2023d) treat input and output functions as graphs, employing message-passing mechanisms to generalize across different domain geometries. However, graph methods are often constrained by local receptive fields, making it difficult to effectively capture global long-range dependencies and suffer from low efficiency on large-scale nodes.
This necessitates the adoption of Transformers (Vaswani et al., 2017). The self-attention mechanism in Transformers is mathematically analogous to learning a global Green’s function (Cao, 2021), making it ideally suited for modeling non-local physical interactions (Alkin et al., 2024; Koshizuka et al., 2024). However, a fundamental scalability barrier remains: the standard self-attention mechanism exhibits quadratic computational complexity $\mathcal{O}(N^{2})$ with respect to the number of mesh points $N$. Given that scientific simulations typically involve $10^{5}$ to $10^{7}$ grid points, standard Transformers are computationally intractable for high-resolution physics. Consequently, the research frontier has shifted towards efficient linear transformers (Li et al., 2023a; Luo et al., 2025; Wu et al., 2024). Recent state-of-the-art methods employ strategies such as low-rank approximations or token clustering to reduce the effective sequence length, achieving linear complexity $\mathcal{O}(N)$ (Rao et al., 2021; Xiao et al., 2024). While these architectures successfully improve throughput, they often do so at the cost of geometric and physical fidelity, as follows:
@@ 2.1 Physics-Informed Neural Networks
Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) represent a foundational paradigm for solving PDEs by embedding governing equations directly into the loss function. PINNs parameterize the solution as a neural network and minimize the PDE residual at collocation points, enabling mesh-free training without labeled simulation data. However, PINNs face fundamental limitations for the problems we target. First, PINNs solve individual PDE instances, requiring complete retraining for each new parameter configuration or boundary condition—a prohibitive cost for design optimization requiring thousands of evaluations. Second, for complex geometries and multi-physics coupling, PINNs frequently encounter ill-conditioning issues (Zhang et al., 2025b, a), often leading to convergence failures or accuracy degradation. Although recent research has alleviated these problems through parameterization (Cho et al., 2024), meta-learning (Cheng and Alkhalifah, 2025), and transfer learning (Wang et al., 2025b) strategies, PINNs still struggle to maintain accuracy and efficiency when the configuration space becomes sufficiently large. These limitations motivate the development of the neural operator paradigm.
@@ 2.2 Neural Operators for PDEs
Neural operators have emerged as a powerful paradigm for learning resolution-invariant mappings between infinite-dimensional function spaces (Kovachki et al., 2023). DeepONet (Lu et al., 2021) pioneered this field by approximating operators through a Branch-Trunk architecture. Subsequently, the Fourier Neural Operator (FNO) (Li et al., 2021) established new benchmarks by parameterizing the integral kernel directly in the frequency domain using FFT. To address the limitations of standard FNO, numerous variants have been proposed: U-NO (Rahman et al., 2023) and U-FNO (Wen et al., 2022) integrate U-Net structures to capture multi-scale features, while F-FNO (Tran et al., 2023) improves computational efficiency by factorizing the Fourier representation across dimensions. To handle irregular geometries beyond rectangular grids, Geo-FNO (Li et al., 2023c) maps physical domains to latent uniform grids, while LSM (Wu et al., 2023) and LNO (Wang and Wang, 2024) apply spectral methods within learned low-dimensional latent spaces to tackle high-dimensional PDE problems. Furthermore, to mitigate the global nature of Fourier transforms that typically leads to loss of local details, wavelet-based methods such as WMT (Gupta et al., 2021) and WNO (Tripura and Chakraborty, 2023) have been introduced to leverage the time-frequency localization properties of wavelets for capturing local variations. For modeling complex unstructured meshes, geometric deep learning offers an alternative perspective. Graph Neural Operators (GNO) (Li et al., 2020) and MeshGraphNet (Pfaff et al., 2021) treat physical domains as graphs, employing message-passing mechanisms to handle complex boundaries. To combine global spectral processing capabilities with local geometric flexibility, hybrid architectures have emerged. For instance, GINO (Li et al., 2023d) integrates Geo-FNO’s global spectral capabilities with GNO’s local graph processing, while 3D-GeoCA (Deng et al., 2024) further incorporates pre-trained 3D vision backbones to enhance geometric feature extraction. However, graph-based methods typically suffer from geometric instability (Morris et al., 2023; Luo et al., 2025) and incur high computational overhead when scaling to million-scale mesh points due to neighborhood search and message passing.
@@ 2.3 Transformers for PDEs
The Transformer architecture has garnered significant attention in scientific computing due to its global receptive field and superior modeling capabilities. Early works such as OFormer (Li et al., 2023a) and HT-Net (Liu et al., 2024) demonstrated the potential of attention mechanisms in capturing non-local physical interactions. However, the quadratic complexity $\mathcal{O}(N^{2})$ of standard attention remains a bottleneck restricting its scalability. To achieve linear complexity $\mathcal{O}(N)$, recent research has explored diverse pathways: AFNO (Guibas et al., 2022), FNet  (Lee-Thorp et al., 2022), and GFNet (Rao et al., 2021) attempt token mixing in the Fourier domain, while SAOT (Zhou et al., 2025) further integrates wavelet attention to mitigate aliasing issues; simultaneously, Galerkin Transformer (Cao, 2021), GNOT (Hao et al., 2023), and ONO (Xiao et al., 2024) utilize kernel tricks or low-rank approximations to achieve linearization; whereas methods represented by FactFormer  (Li et al., 2023b) and Transolver (Wu et al., 2024) employ axial decomposition or token aggregation strategies, respectively. Despite their significant efficiency improvements, these dimensionality reduction or aggregation operations act as implicit spatial low-pass filters, inevitably discarding high-frequency geometric details. Furthermore, they typically allocate a uniform computational budget across the entire domain, failing to adapt to physical heterogeneity in flow fields.
@@ 4.1 Experimental Setup
Our evaluation encompasses six benchmarks with varying geometric complexities. The standard benchmarks (Li et al., 2021, 2023c) include Elasticity (inner stress prediction on 2D point clouds), Plasticity (temporal displacement forecasting under external forces), Airfoil (transonic flow simulation with shock waves), and Pipe (velocity field estimation in curved structures). For industrial-scale validation, we employ Shape-Net Car (Umetani and Bickel, 2018) for automotive aerodynamic simulation on hybrid 3D geometries, and AirfRANS (Bonnet et al., 2022) for high-fidelity RANS simulation on NACA airfoils under varying Reynolds numbers and angles of attack. Detailed dataset descriptions are provided in  Appendix A.1.
@@
We compare against representative methods from three categories: spectral neural operators (FNO (Li et al., 2021), Geo-FNO (Li et al., 2023c), U-NO (Rahman et al., 2023), F-FNO (Tran et al., 2023), LSM (Wu et al., 2023), WMT (Gupta et al., 2021)); Transformer-based approaches (Galerkin Transformer (Cao, 2021), OFormer (Li et al., 2023a), GNOT (Hao et al., 2023), ONO (Xiao et al., 2024), FactFormer (Li et al., 2023b), IPOT (Lee and Oh, 2024), Transolver (Wu et al., 2024), SAOT (Zhou et al., 2025)); and geometric deep learning methods (PointNet (Qi et al., 2017), Graph U-Net (Gao and Ji, 2019), GraphSAGE (Hamilton et al., 2017), MeshGraphNet (Pfaff et al., 2021), GNO (Li et al., 2020), GINO (Li et al., 2023d), 3D-GeoCA (Deng et al., 2024)). See  Appendix A.4 for details.
@@
We adopt the relative $L^{2}$ error as the primary evaluation metric. For industrial benchmarks, we additionally report aerodynamic coefficient errors and Spearman’s rank correlation. PGOT is configured with $L=8$ PhysGeoBlocks, with other hyperparameters adapted to each task. All experiments are conducted on NVIDIA A100 GPUs and repeated three times. Comprehensive details on metrics, training configurations, and hyperparameters are provided in  Appendix A.2 and  Appendix A.3.
@@ 4.2 Main Results
Table 2 summarizes results on large-scale industrial benchmarks. On Shape-Net Car, PGOT achieves the best performance across all metrics, with 9.1% reduction in volumetric field error and 10.6% improvement in drag coefficient prediction compared to Transolver. The Spearman correlation $\rho_{D}=0.9926$ indicates excellent ranking capability for design optimization. Figure 5(b) visualizes these results, showing excellent agreement between PGOT’s predicted streamlines and ground truth, with superior accuracy in geometrically complex regions like wheel wells. On AirfRANS, PGOT demonstrates substantial improvements: 32.4% reduction in volumetric field error, 81.3% reduction in surface field error, and 12.2% improvement in lift coefficient prediction. As shown in Figure 5(a), while Transolver and SAOT produce radial error patterns, PGOT achieves significantly cleaner predictions with errors primarily concentrated in the far-field region. These results underscore PGOT’s capability in handling complex unstructured meshes with over 32,000 points while maintaining high prediction fidelity. Efficiency analysis is detailed in Appendix B.2.
@@ 4.3 Ablation Studies
Table 4 investigates the effect of multi-scale geometric encoding. Single-scale encoding ($S=1$) performs poorly across all benchmarks, confirming that aggregation-induced geometric aliasing significantly degrades prediction quality. Performance improves substantially and reaches optimality at $S=2$, demonstrating that a dual-scale representation is sufficient to capture both global topology and local boundary details. Interestingly, further increasing the scales to $S=3$ or $S=4$ leads to performance degradation. This suggests that excessive high-frequency geometric priors may introduce spectral noise or redundancy, which interferes with the model’s ability to learn robust physical dynamics. See  Appendix C for additional visualizations.
@@ 5 Conclusion
Appendix Appendix A Implementation Details
Appendix A.1 Dataset Descriptions
@@
This industrial benchmark focuses on automotive aerodynamic simulation at 72 km/h driving conditions. The task requires simultaneous prediction of surface pressure and surrounding air velocity. Vehicle geometries are sourced from the “car” category of the ShapeNet repository and discretized into unstructured meshes with approximately 32,000 points combining surface and volumetric elements. The input is preprocessed to include mesh point positions, signed distance functions, and normal vectors. The predicted fields can be used to compute drag coefficients for design optimization. The training set contains 789 samples, and the test set contains 100 samples (Umetani and Bickel, 2018).
@@ Appendix A.2 Evaluation Metrics
Appendix A.3 Training and Model Configurations
Following the conventions established in prior work (Li et al., 2021, 2023c; Wu et al., 2024; Zhou et al., 2025), our model is trained using the relative $L^{2}$ loss. Table 8 summarizes the training configurations shared across all methods for fair comparison, and Table 9 presents PGOT-specific hyperparameters.
@@ Appendix A.3 Training and Model Configurations
For Shape-Net Car (Umetani and Bickel, 2018) and AirfRANS (Bonnet et al., 2022), $\mathcal{L}_{v}$ and $\mathcal{L}_{s}$ denote the relative $L^{2}$ losses on volumetric and surface fields.
@@ Appendix A.4 Baseline Implementations
For FNO (Li et al., 2021), Geo-FNO (Li et al., 2023c), WMT (Gupta et al., 2021), U-FNO (Wen et al., 2022), U-NO (Rahman et al., 2023), F-FNO (Tran et al., 2023), and LSM (Wu et al., 2023), we report results from the original publications (Wu et al., 2024; Zhou et al., 2025; Li et al., 2021).
@@
For Galerkin Transformer (Cao, 2021), OFormer (Li et al., 2023a), GNOT (Hao et al., 2023), HT-Net (Liu et al., 2024), FactFormer (Li et al., 2023b), and ONO, we report results from the original publications (Wu et al., 2024; Zhou et al., 2025; Li et al., 2021). For Transolver (Wu et al., 2024) and SAOT (Zhou et al., 2025), we reproduce results based on their official codebases with all hyperparameter configurations kept consistent with PGOT to ensure fair comparison.
@@
For MLP, GraphSAGE (Hamilton et al., 2017), PointNet (Qi et al., 2017), Graph U-Net (Gao and Ji, 2019), MeshGraphNet (Pfaff et al., 2021), GNO (Li et al., 2020), GINO (Li et al., 2023d), and 3D-GeoCA (Deng et al., 2024), we report results from the original publications.
Appendix Appendix B Supplementary Results
Appendix B.1 Ablation Results